# Fix-Con: Automatic Fault Localization and Repair of Deep Learning Model Conversions between Frameworks

### v3

@@ Abstract
In this paper we propose an automated approach for fault localization and repair, Fix-Con, during model conversion between deep learning frameworks. Fix-Con is capable of detecting and fixing faults introduced in model input, parameters, hyperparameters, and the model graph during conversion.
@@
Fix-Con uses a set of fault types (mined from surveying conversion issues reported in code repositories and forums) to localize potential conversion faults in the converted target model and then repair them appropriately, e.g., replacing the parameters of the target model with those from the source model. This is done iteratively for every image in the dataset, comparing output label differences between the source model and the converted target model until all differences are resolved. We evaluate the effectiveness of Fix-Con in fixing model conversion bugs of three widely used image recognition models converted across four different deep learning frameworks. Overall, Fix-Con was able to fix $462462$ out of $755755$ detected conversion faults, either completely repairing or significantly improving the performance of $1414$ out of the $1515$ erroneous conversion cases.
I Introduction
@@
With the widespread use of deep learning (DL) in various domains, there is an inherent need for DL models to be inter-operable across DL frameworks (such as PyTorch [67], TensorFlow (TF) [64], Keras [54]) to maximize re-usability of models across frameworks. Conversion of DL models between frameworks is facilitated by a plethora of automated conversion tools such as tf2onnx [44], onnx2keras [45], onnx2torch [46], MMdnn [61], among others. However, the conversion process may be riddled with bugs [63, 62, 43], making the converted models undeployable, perform poorly in terms of output label correctness when changing from one framework to another, run slowly, or face challenges in robust deployment [66, 52, 53].
@@
Louloudakis et al. [63] provided empirical evidence of model conversion errors and resulting output label errors in four DL frameworks. In this paper, we target the conversion errors they observed for fault localization and repair. They considered three image recognition models pre-trained on ImageNet [55] that were treated as Source models. They then converted each Source model to use a different DL framework, referred to as Target. Figure 1 from [63] shows the proportion of output label dissimilarities between (Source, Target) pairs across all images in the test dataset. The conversion process failed to execute in $1111$ out of the $3636$ conversions, seen as empty gray cells in Figure 1. For the remaining conversions, they found varying levels of image label discrepancies between Target and Source models from 0 to 100%, with $1515$ cases having non-zero percentage of output label dissimilarity. This study underscores the error-prone nature of the conversion process, highlighting the necessity for a technique to localize and repair faults introduced by DL framework converters.
@@
Figure 1: Pairwise comparison of output labels between Source and converted Target models (from [63]).
@@
To address this problem, in this paper we propose Fix-Con, an automated approach for fault localization and repair of erroneous model conversions between DL frameworks. We focus on the error cases in Figure 1 with non-zero output label differences. The gray boxes in Figure 1, where the conversion tools crashed, have been explored by existing tools like MMdnn [61] and complements the capability we provide in diagnosing and repairing conversions with output label differences.
@@
Fix-Con is capable of detecting and fixing faults introduced in model layer weights, biases, hyperparameters, and the model computation graph during conversion. It compares the Source model (in one DL framework) against the converted Target model (in a different framework) using an input dataset. For images with label discrepancies between Source and Target, Fix-Con assesses the difference between the models with respect to their parameters. Then it uses a set of strategies to mitigate conversion errors such as the replacement of Target model parameters with those from Source.
@@
Our approach for Fix-Con focuses on the Target model rather than the converter tool itself for a variety of reasons. First, errors observed in conversions may be related to problematic configurations (e.g., input preprocessing) that are not related to the conversion tool. Second, the source code of the conversion tool may be unavailable or proprietary, and therefore not accessible for direct repair. Third, the conversion process might involve manual intervention or conversion (e.g., [25, 6]), which in turn can introduce errors. Fourth, our methodology is not tied to any specific conversion tool/version, and can be applied to any model conversion between source and target settings, including intermediate representations like ONNX [47]. It is worth noting that conversion tools are constantly evolving, thus having a localize/repair approach independent of the tool source code accessibility and evolution would prove helpful and more dependable for AI model users. As an aside, the fault localization and repair techniques we propose can also be adopted by conversion tool developers to address faults associated with the tools themselves.
@@
We evaluate the effectiveness of Fix-Con in fixing the model conversion errors reported by Louloudakis et al [63]. Fix-Con was able to effectively localize and repair $1414$ of the $1515$ erroneous model conversions that had non-zero output label dissimilarities.
@@
    An automated fault repair framework, Fix-Con, that can fix converted Target models to match output labels with that of the Source model.
  3. 3.
@@
    A comprehensive empirical evaluation of the effectiveness of Fix-Con in fault localization and repair.
II Background
@@
We provide a brief background on deep learning frameworks, conversions between them, the intermediate ONNX format and the compiler framework Apache TVM that we use for implementation and analysis with Fix-Con.
II-A Deep Learning Frameworks
@@
Deep Learning Frameworks provide utilities to machine learning engineers and developers in order to build, train, deploy and optimize Deep Neural Networks (DNNs). We focus on four actively maintained and widely-used DL frameworks in our experiments: Keras [54], PyTorch [67], TensorFlow (TF) [64], and TensorFlow Lite (TFLite - a framework focused on edge devices) [64].
II-B Framework Conversions
@@
DNN models are converted from one DL framework to another, for two primary reasons: (1) portability - so that a model can be deployed on devices of different computational capabilities; and (2) supported features and/or operations, such as extended optimization support or better profiling options. The conversion process is conducted by providing a pre-trained model built on one DL framework (e.g., PyTorch). We call this model Source. Then, the conversion tool will use Source in order to generate the same model, but built in a new DL framework (e.g., Keras). We call this model Target. There is a wide variety of open-source tools for this purpose [44, 45, 46, 61], and the conversion procedure is performed automatically. However, the conversion process might involve multiple converters and might also require the model to be built using an intermediate representation (e.g., ONNX [47]). For instance, the conversion of a model from PyTorch to Keras might require a two-step conversion: (1) PyTorch to ONNX, using the native converter of PyTorch to generate ONNX; and (2), the onnx2keras [45] converter, in order to convert the ONNX intermediate representation to Keras.
II-C Open Neural Network Exchange (ONNX)
@@
ONNX [47] is an open standard for machine learning interoperability that allows the representation of DNNs using a common set of operators and file formats to enable interoperability with a variety of tools, and systems. ONNX provides a powerful API in order to analyze model graph, hyperparameters, and parameters. The model is represented by a graph of operations (or nodes), each containing a set of properties. These operations are executed sequentially, propagating data from one layer to its subsequent ones. Throughout the years, ONNX evolves producing newer versions of the standard, modified from the past. Each standard version is attributed an opset number. Due to the popularity of ONNX and the versatility of its API, we perform our automatic fault localization and repair in this representation level - retrieving ONNX versions of the Source and Target model, analyzing them and repairing them. For that purpose, when we refer to Source and Target  we implicitly refer to the ONNX variants of the models under test.
II-D Apache TVM
@@
Apache TVM [51] is an end-to-end machine learning compiler framework for CPUs, GPUs, and specialized accelerators. It is actively used and supported by a wide community of developers and researchers worldwide. TVM allows compilation, optimization and deployment of DNNs in a variety of hardware acceleration devices. It also features a very powerful debugger, which allows to export metadata regarding the model inference, such as layer activations, execution times per-model layer, as well as metadata about model structure and parameters, each exported in separate metadata files, while also providing a high-level API for their analysis. We utilize the TVM Debugger in order to perform Layer Activation Analysis, as described in Section IV-A4.
III Related Work
@@
In terms of model conversions, limited research is conducted in the literature. Openja et al. [65] published a study that emphasizes the challenges of the conversion process. Louloudakis et al. [63, 62] explored the effects of the computational environment parameters on image recognition models, including deep learning (DL) framework conversions. Liu et al. [61] propose MMdnn, a framework to automate the framework conversion process and mitigate a number of conversion challenges that result in conversion failures, such as unavailable operators, inconsistent tensor layouts, unsupported padding and incompatible argument types. MMdnn, however, does not consider faults related to configuration, such as input preprocessing, or faults related to additional important aspects such as layer parameters and computation graph that can also result in output incompatibilities during conversion. This is addressed by Fix-Con. It is also worth noting that MMdnn is currently in a deprecated state, with its last update on August 2020. In addition, our efforts with testing MMdnn was unsuccessful as it was built with older, deprecated versions of DL frameworks (e.g., PyTorch v0.4.0, while the current version is v2.2.0).
@@
In an attempt to compare MMdnn with our work, we identified a potential overlap of only $1515$ potential faults out of the $462462$ we repaired with Fix-Con, based on the results reported in [61]. Finally, MMdnn supports conversions over a restricted set of DL frameworks (primarily CNTK, CoreML, Keras, MXNet, PyTorch, TensorFlow, Caffe, and DarkNet) that is not currently maintained or updated while our approach is generalizable to model conversions between any pair of DL frameworks.
@@
Automatic fault localization of DNNs is also explored in the literature DeepFault [56] focuses on a suspiciousness-oriented spectrum analysis algorithm to identify problematic parts of the model that can be responsible for faults, and also proposes a method for adversarial input generation. DeepLocalize [74] converts the model into an imperative representation and then performs dynamic analysis on its execution traces. In the context of fault localization of DL frameworks, CRADLE [68] performs an execution analysis on the model graph in order to detect faults, while LEMON [73] utilizes the metrics used by CRADLE and applies mutation testing in order to identify issues in the model. However, none of the aforementioned approaches focus on faults encountered during model conversions.
To achieve model conversions between DL frameworks, a variety of popular tools exist, such as tf2onnx [44], onnx2keras [45], onnx2torch [46], tflite2onnx [42], and native DL framework converters such as TFLiteConverter from TFLite [64], and ONNX converter from PyTorch [67]. However, these tools come with no guarantees, and can potentially introduce faults in the converted models.
IV Methodology
@@
Figure 2: Fault localization and repair pipeline of Fix-Con handling 6 fault types, three within Input-based category and three within Layer-based category.
@@
Fix-Con first takes as input the dataset, Source and Target models. It then localizes and repairs conversion errors that resulted in output label discrepancies between Source and Target. To compare the Source and Target model architectures, configuration and parameters we use their corresponding representation in the ONNX format which is a popular intermediate representation. We then import the Source-ONNX and Target-ONNX models to the TVM compiler framework to facilitate comparison. Fix-Con comprises two main components: 1) A Fault Localization component that identifies locations, parameters and properties in the Target model that may be potentially faulty; and 2) a Fault Repair component with a defined set of strategies to fix the faulty locations and aspects identified through fault localization. The full fault localization and repair pipeline of Fix-Con is presented in Figure 2.
IV-A Fault Localization
@@
We surveyed $4040$ posts and issues from StackOverflow, GitHub, and forums such as TensorFlow and NVidia discussions, where model conversion errors were raised, to understand the possible locations of errors. Table I shows six different conversion fault types identified in these issues that were related to 1) Preprocessing (PP), (2) Input Dimensions (ID), (3) Tensor Shape and Structure (TSS), (4) Weights & Biases (WB), (5) Layer Hyperparameters (LH), and (6) Computation Graph (CG). We also discovered many of these fault types examining the conversion errors reported in [63].
@@
TABLE I: Model conversion faults from StackOverflow, GitHub issues and other forums. Faults are classified as input-based and layer-based.
@@
1 Input $P$ [3, 19, 10, 39, 7, 30]
2 Input ID [5, 36, 37, 32, 13, 20, 23, 26, 28]
3 Input TSS [11, 14, 21, 27, 31, 41]
4 Layer WB [8, 9, 12, 16, 2, 29, 34, 1, 15, 35, 33, 40]
5 Layer LH [17, 18, 50]
6 Layer CG [4, 38, 22, 24]
As seen in Figure 2, our fault localization approach starts by localizing the first three fault types: PP, ID, and TSS. These three fault types pertain to the input, stemming from differences between Source and Target with respect to the preprocessing used, or changes in input dimensions or operations on the input. We refer to these fault types as Input-based. After attempting to localize input-based fault types, our approach proceeds with the remaining three fault types in Table I that belong to the Layer-based category, as it stems from differences in layer hyperparameters, weights and biases, and differences in computation graph.

### v4

# FetaFix: Automatic Fault Localization and Repair of Deep Learning Model Conversions

@@ Abstract.
In this paper, we propose an automated approach for fault localization and repair, FetaFix, during model conversion between deep learning frameworks. FetaFix is capable of detecting and fixing faults introduced in model input, parameters, hyperparameters, and the model graph during conversion. FetaFix uses a set of fault types (mined from surveying common conversion issues reported in code repositories and forums) to localize potential conversion faults in the converted target model and then repair them appropriately, e.g., replacing the parameters of the target model with those from the source model. This is done iteratively for every image in the dataset, comparing output label differences between the source model and the converted target model until all differences are resolved. We evaluate the effectiveness of FetaFix in fixing model conversion bugs of three widely used image recognition models converted across four different deep learning frameworks. Overall, FetaFix was able to fix $462462$ out of $755755$ detected conversion faults, either completely repairing or significantly improving the performance of $1414$ out of the $1515$ erroneous conversion cases.
conference: The 29th International Conference on Evaluation and Assessment in Software Engineering; 17–20 June, 2025; Istanbul, Türkiyepublicationmonth: 06journalyear: 2025
1. Introduction
@@
With the widespread use of deep learning (DL) in various domains, there is an inherent need for DL models to be inter-operable across DL frameworks (such as PyTorch (Paszke et al., 2019), TensorFlow (TF) (Martín Abadi et al., 2015), Keras (Chollet, 2015)) to maximize re-usability of models across frameworks. Conversion of DL models between frameworks is facilitated by a plethora of automated conversion tools such as tf2onnx (tf2, 2022), onnx2keras (onn, 2023a), onnx2torch (onn, 2023b), MMdnn (Liu et al., 2020), among others. However, the conversion process may be riddled with bugs (Louloudakis et al., 2023a, 2022; TF-, 2021), making the converted models undeployable, perform poorly in terms of output label correctness when changing from one framework to another, run slowly, or face challenges in robust deployment (Openja et al., 2022b; Chen et al., 2020, 2021).
@@
In a previous study of ours (Louloudakis et al., 2023a), we showcased empirical evidence of problems in model conversion across four well-known DL frameworks. In particular, we considered three image recognition models, pre-trained on ImageNet (Deng et al., 2009), that were treated as Source models. We then converted each Source model to use a different DL framework, referred to as Target. Figure 1 from (Louloudakis et al., 2023a) shows the proportion of output label dissimilarities between (Source, Target) pairs across all images in the test dataset. The conversion process failed to execute in $1111$ out of the $3636$ conversions, seen as empty gray cells in Figure 1. For the remaining conversions, they found varying levels of image label discrepancies between Target and Source models from 0 to 100%, with $1515$ cases having non-zero percentage of output label dissimilarity. This study underscores the error-prone nature of the conversion process, highlighting the necessity for a technique to localize and repair faults introduced by DL framework converters.
@@
Figure 1. Pairwise comparison of output labels between Source and converted Target models (from (Louloudakis et al., 2023a)).
@@
To address this problem, in this paper we propose FetaFix [^1] The source code (accompanied with detailed instructions), as well a small dataset for evaluation are available at https://github.com/luludak/FetaFix, an automated approach for fault localization and repair of erroneous model conversions between DL frameworks. We focus on the cases in Figure 1 that did not crash, but resulted in non-zero output label differences across Source and Target. The gray boxes in Figure 1, where the conversion tools crashed, have been explored by existing tools like MMdnn (Liu et al., 2020) and complements the capability we provide in diagnosing and repairing conversions with output label differences.
@@
FetaFix is capable of detecting and fixing faults introduced in model layer weights, biases, hyperparameters, and the model computation graph during conversion. It compares the Source model (in one DL framework) against the converted Target model (in a different framework) using an input dataset. For images with label discrepancies between Source and Target, FetaFix assesses the difference between the models with respect to their parameters. Then it uses a set of strategies to mitigate conversion errors such as the replacement of Target model parameters with those from Source.
@@
Our approach for FetaFix focuses on the Target model rather than the converter tool itself for a variety of reasons. First, errors observed in conversions may be related to problematic configurations (e.g., input preprocessing) that are not related to the conversion tool. Second, the source code of the conversion tool may be unavailable or proprietary, and therefore not accessible for direct repair. Third, the conversion process might involve manual intervention or conversion (e.g., (man, 2024b, a)), which in turn can introduce errors. Fourth, our methodology is not tied to any specific conversion tool/version, and can be applied to any model conversion between source and target settings, including intermediate representations like ONNX (onn, 2023c). It is worth noting that conversion tools are constantly evolving, thus having a localize/repair approach independent of the tool source code accessibility and evolution would prove helpful and more dependable for AI model users. As an aside, the fault localization and repair techniques we propose can also be adopted by conversion tool developers to address faults associated with the tools themselves.
@@
Focusing on the non-crashing conversions from our prior work (Louloudakis et al., 2023a), we evaluated the effectiveness of FetaFix in fixing model conversion errors. FetaFix was able to effectively localize and repair $1414$ of the $1515$ erroneous model conversions that had non-zero output label dissimilarities.
@@
    An automated fault repair framework, FetaFix, that can fix converted Target models to match output labels with that of the Source model.
  3. (3)
@@
    A comprehensive empirical evaluation of the effectiveness of FetaFix in fault localization and repair.
2. Background
@@
We provide a brief background on deep learning frameworks, conversions between them, the intermediate ONNX format and the compiler framework Apache TVM that we use for implementation and analysis with FetaFix.
2.1. Deep Learning Frameworks
@@
Deep Learning Frameworks provide utilities to machine learning engineers and developers in order to build, train, deploy and optimize Deep Neural Networks (DNNs). We focus on four actively maintained and widely-used DL frameworks in our experiments: Keras (Chollet, 2015), PyTorch (Paszke et al., 2019), TensorFlow (TF) (Martín Abadi et al., 2015), and TensorFlow Lite (TFLite - a framework focused on edge devices) (Martín Abadi et al., 2015).
2.2. Framework Conversions
@@
DNN models are converted from one DL framework to another, for two primary reasons: (1) portability - so that a model can be deployed on devices of different computational capabilities; and (2) supported features and/or operations across the Deep Learning Acceleration Stack (Gibson et al., 2025), such as extended optimization support or better profiling options. The conversion process is conducted by providing a pre-trained model built on one DL framework (e.g., PyTorch). We call this model Source. Then, the conversion tool will use Source in order to generate the same model, but built in a new DL framework (e.g., Keras). We call this model Target. There is a wide variety of open-source tools for this purpose (tf2, 2022; onn, 2023a, b; Liu et al., 2020), and the conversion procedure is performed automatically. However, the conversion process might involve multiple converters and might also require the model to be built using an intermediate representation (e.g., ONNX (onn, 2023c)). For instance, the conversion of a model from PyTorch to Keras might require a two-step conversion: (1) PyTorch to ONNX, using the native converter of PyTorch to generate ONNX; and (2), the onnx2keras (onn, 2023a) converter, in order to convert the ONNX intermediate representation to Keras.
2.3. Open Neural Network Exchange (ONNX)
@@
ONNX (onn, 2023c) is an open standard for machine learning interoperability that allows the representation of DNNs using a common set of operators and file formats to enable interoperability with a variety of tools, and systems. ONNX provides a powerful API in order to analyze model graph, hyperparameters, and parameters. The model is represented by a graph of operations (or nodes), each containing a set of properties. These operations are executed sequentially, propagating data from one layer to its subsequent ones. Throughout the years, ONNX evolves producing newer versions of the standard, modified from the past. Each standard version is attributed an opset number. Due to the popularity of ONNX and the versatility of its API, we perform our automatic fault localization and repair in this representation level - retrieving ONNX versions of the Source and Target model, analyzing them and repairing them. For that purpose, when we refer to Source and Target  we implicitly refer to the ONNX variants of the models under test.
2.4. Apache TVM
@@
Apache TVM (Chen et al., 2018) is an end-to-end machine learning compiler framework for CPUs, GPUs, and specialized accelerators. It is actively used and supported by a wide community of developers and researchers worldwide. TVM allows compilation, optimization and deployment of DNNs in a variety of hardware acceleration devices. It also features a very powerful debugger, which allows exporting metadata regarding the model inference, such as layer activations, execution times per-model layer, as well as metadata about model structure and parameters, each exported in separate metadata files, while also providing a high-level API for their analysis. We utilize the TVM Debugger in order to perform Layer Activation Analysis, as described in Section 4.1.4.
3. Related Work
@@
To achieve model conversions between DL frameworks, a variety of tools exist, such as tf2onnx (tf2, 2022), tflite2onnx (tfl, 2020), onnx2keras (onn, 2023a), onnx2torch (onn, 2023b), and native DL framework converters such as ONNX converter from PyTorch (Paszke et al., 2019) and TFLiteConverter from TFLite (Martín Abadi et al., 2015). However, these tools come with no guarantees, and can potentially introduce faults in the converted models. To this direction however, limited research is conducted in the literature. Openja et al. (Openja et al., 2022a) published a study that emphasizes the challenges of the conversion process. Jajal et al. (Jajal et al., 2024) conducted a failure analysis on ONNX model converters, concluding that converters are responsible for about 75% of the conversion defects, and that 33% of reported failures are related to semantically incorrect models. Liu et al. (Liu et al., 2020) propose MMdnn, a framework to automate the framework conversion process and mitigate a number of conversion challenges that result in conversion failures, such as unavailable operators, inconsistent tensor layouts, unsupported padding and incompatible argument types. MMdnn, however, does not consider faults related to configuration, such as input preprocessing, or faults related to additional important aspects such as layer parameters and computation graph that can also result in output incompatibilities during conversion. This is addressed by FetaFix. It is also worth noting that MMdnn is currently in a deprecated state, with its last update on August 2020. In addition, our efforts with testing MMdnn was unsuccessful as it was built with older, deprecated versions of DL frameworks (e.g., PyTorch v0.4.0, while the current version is v2.2.0). We rigorously (Louloudakis et al., 2023b, 2022, a) explored the effects of the computational environment parameters on image recognition models, including deep learning (DL) framework conversions, and implemented a proof-of-concept (Louloudakis et al., 2024) concerning conversion faults related to model weights and biases, which is the predecessor of FetaFix.
@@
In an attempt to compare MMdnn with our work, we identified a potential overlap of only $1515$ potential faults out of the $462462$ we repaired with FetaFix, based on the results reported in (Liu et al., 2020). Finally, MMdnn supports conversions over a restricted set of DL frameworks (primarily CNTK, CoreML, Keras, MXNet, PyTorch, TensorFlow, Caffe, and DarkNet) that is not currently maintained or updated while our approach is generalizable to model conversions between any pair of DL frameworks.
@@
Automatic fault localization of DNNs is also explored in the literature. DeepFault (Eniser et al., 2019) focuses on a suspiciousness-oriented spectrum analysis algorithm to identify problematic model parts, and proposes a method for adversarial input generation. DeepLocalize (Wardat et al., 2021) converts the model into an imperative representation and then performs dynamic analysis on its execution traces. In the context of fault localization of DL frameworks, CRADLE (Pham et al., 2019) performs an execution analysis on the model graph in order to detect faults, while LEMON (Wang et al., 2020) utilizes the metrics used by CRADLE and applies mutation testing to identify issues in the model. However, none of the aforementioned approaches focus on faults encountered during model conversions.
4. Methodology
@@
Figure 2. Fault localization and repair pipeline of FetaFix handling 6 fault types, three within Input-based category and three within Layer-based category.
@@
FetaFix first takes as input the dataset, Source and Target models. It supports any image test dataset, requiring only the definition of a dataset path in its configuration. It then localizes and repairs conversion errors that resulted in output label discrepancies between Source and Target. To compare the Source and Target model architectures, configuration and parameters we use their corresponding representation in the ONNX format which is a popular intermediate representation. We then import the Source-ONNX and Target-ONNX models to the TVM compiler framework to facilitate comparison. FetaFix comprises two main components: 1) A Fault Localization component that identifies locations, parameters and properties in the Target model that may be potentially faulty; and 2) a Fault Repair component with a defined set of strategies to fix the faulty locations and aspects identified through fault localization. The full fault localization and repair pipeline of FetaFix is presented in Figure 2.
4.1. Fault Localization
@@
We surveyed $4040$ posts and issues from StackOverflow, GitHub, and forums such as TensorFlow and NVIDIA discussions, where model conversion errors were raised, to understand the possible locations of errors. Although not exhaustive, we believe this sample to be representative of issues faced by developers, as (1) the issues we observed were among the most popular with regard to conversions and (2) the same or very similar issues appeared across different trackers and forums. Table 1 shows six different conversion fault types identified in these issues that were related to 1) Preprocessing (PP), (2) Input Dimensions (ID), (3) Tensor Shape and Structure (TSS), (4) Weights & Biases (WB), (5) Layer Hyperparameters (LH), and (6) Computation Graph (CG). We also discovered many of these fault types examining the conversion errors reported in (Louloudakis et al., 2023a).
@@
Table 1. Model conversion faults from StackOverflow, GitHub issues and other forums. Faults are classified as input-based and layer-based.
@@
1 Input $P$ (sta, 2023b, f, d, n; pyt, 2023; ten, 2024)
2 Input ID (sta, 2023c; ten, 2023c; sta, 2023m, k, e; git, 2023j; sta, 2023g; git, 2023m; sta, 2023i)
3 Input TSS (git, 2023d, f, k, n, o, t)
4 Layer WB (git, 2023b, c, e, g; ten, 2023a; sta, 2023j, l, a; ten, 2023b; git, 2023q, p, s)
5 Layer LH (git, 2023h, i; ACervantes, 2021)
6 Layer CG (git, 2023a, r, l; sta, 2023h)
As seen in Figure 2, our fault localization approach starts by localizing the first three fault types: PP, ID, and TSS. These three fault types pertain to the input, stemming from differences between Source and Target with respect to the preprocessing used, or changes in input dimensions or operations on the input. We refer to these fault types as Input-based. After attempting to localize input-based fault types, our approach proceeds with the remaining three fault types in Table 1 that belong to the Layer-based category, as it stems from differences in layer hyperparameters, weights and biases, and differences in computation graph.